OpenStudy (anonymous):

Formulas for 6? (I have all the other formulas, for 1 2 3 4 and 5, in case anyone needs)

OpenStudy (anonymous):

I have the following values :V0=10V, T=0.5ms, Rw=1Ω, Lw=0.5mH and RL=4Ω.

OpenStudy (anonymous):

Q2 P3 formula benzo is correct. How are you inputing your values? Give them to me and let me compute, then we compare.

OpenStudy (anonymous):

did it give 2.60944 to you?

OpenStudy (anonymous):

@ Jone133, we have the same values. a=0.1, b=0.1, c= 7.942. still having problems with d.

OpenStudy (anonymous):

@kidarth what about 1.69315?

OpenStudy (anonymous):

yeah i Checked it its right given your values

OpenStudy (anonymous):

i got mine right in 2nd try

OpenStudy (anonymous):

V0=10V, T=0.5ms, Rw=1Ω, Lw=0.5mH and RL=4Ω.

OpenStudy (anonymous):

i said i checked your answer by changing my values to yours just after i completed my analysis so i think the anser 1.69 is right

OpenStudy (anonymous):

@kidarth i said i checked your answer by changing my values to yours just after i completed my analysis so i think the anser 1.69 is right Happy 100%

OpenStudy (anonymous):

gooood I think i need government authorised documents and signatures to make you believe me i didnt use any formula but instead used transient analysis if you think my answer is right, use it ,rest is your choice

OpenStudy (anonymous):

Q2: q2 p1: Vs as vin<vt p2: v0 = vs-kr(e^((vin-vt)/vn)-1) p3: vin = ln(vs/2kr +1)vn+vt p4: vin=sqrt(vs/kr)+vt p5: -kre^((vin-vt)/vn) p6: -kr(vin-vt) vs=4 r =1k k = 1 VT = 1 Vn = 1 vs=4.5 r = 1k k =1mA vt=1 vn=1 1: 4.5 vin=0.5 2: 2.7818 vin=2 3: 2.1786 vo=vs/2 4: 3.1213 vo=vs/2 mosfet 5: -2.7182 vin=2 6: -1 vin=2 mosfet

OpenStudy (anonymous):

formula for question 1?